Show that the polynomial 2x + 1 in Z4[x] has a multiplicative inverse in Z4[x].


1
Expert's answer
2022-01-04T17:55:11-0500

Solution:

Assume "(1+2x).(1+2x)"

"(1 + 2x)^2 = 1 + 4x + 4x^2 = 1."

The inverse of "2x + 1" in "Z_4[x]" is "2x + 1".

Thus, the polynomial "2x + 1" in "Z_4[x]" has a multiplicative inverse in "Z_4[x]".